OpenStudy (mallorysipp234):

Given the parent functions f(x) = 5x − 1 and g(x) = 3x − 9, what is g(x) − f(x)?

OpenStudy (mallorysipp234):

g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x − 10 − 5x g(x) − f(x) = −2x − 8 − 3x g(x) − f(x) = −2x − 10 − 3x

OpenStudy (mallorysipp234):

and could you please explain, i dont understand.

jigglypuff314 (jigglypuff314):

plug in f(x)=5x-1 for f(x) and g(x)=3x-9 for g(x) so g(x) - (fx) = (3x-9) - (5x-1) simplify

OpenStudy (mallorysipp234):

so out of (3x-9) - (5x-1) would i combine the like terms first?

jigglypuff314 (jigglypuff314):

yes :) but first I suggest you distribute the negative of -(5x-1)

OpenStudy (mallorysipp234):

okay so -5x + 1

OpenStudy (mallorysipp234):

would it be letter C? :)

jigglypuff314 (jigglypuff314):

mmm not quite :) after distributing you get 3x - 9 - 5x + 1

OpenStudy (mallorysipp234):

so then it would be A, after you add the -9 to the 1, right? making that a -8.

jigglypuff314 (jigglypuff314):

yep :) that's what I got ^.^

OpenStudy (mallorysipp234):

can you help me with a few more? its not much. Just 3. please!

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (mallorysipp234):

Given the function f(x) = 0.3(4)x, what is the value of f−1(6)? 0.360 1.242 2.161 2.312

jigglypuff314 (jigglypuff314):

oh... is that an inverse?

OpenStudy (mallorysipp234):

to be honest i have no idea.

jigglypuff314 (jigglypuff314):

I'm confused on this one sorry :/